# Diagnosing crosstalk in large-scale QPUs using zero-entropy classical shadows

### v2

@@ Abstract
As quantum processing units (QPUs) scale toward hundreds of qubits, diagnosing crosstalk and noise-induced correlations becomes critical for reliable quantum computation. In this work, we introduce Zero-Entropy Classical Shadows (ZECS), a diagnostic tool that reconstructs positive semidefinite, unit-trace density operators from classical shadow (CS) information. ZECS enables proper subregion tomography and reduces the effect of sampling and time-dependent errors. We apply ZECS to large superconducting QPUs, including ibm_brisbane (127 qubits) and ibm_fez (156 qubits), using 6,000 samples. With these samples, ZECS detects and characterizes crosstalk among disjoint qubit subsets across the full hardware topology. This information is then used to select low-crosstalk qubit subsets for executing the Quantum Approximate Optimization Algorithm (QAOA) on a 20-qubit problem. Compared to the best qubit selection via Qiskit transpilation, our method improves solution quality by 10% and increases algorithmic coherence by 33%. ZECS offers a scalable and measurement-efficient approach to diagnosing noise in large-scale QPUs.
@@
zero-entropy classical shadow, quantum state tomography, routing qubits, IBM Quantum, QPU crosstalk.
@@ I Introduction
Figure 1: (a) Shows the ZECS density state operator reconstruction workflow. (b) Shows the QPU layout with three sets of qubits chosen to implement some circuits. Vertices represent qubits and edges two-qubit native interactions of the QPU. (c) Shows the acquisition of the CS information to reconstruct the density state operator. After implementing the circuit, the measuring basis changes randomly from the Pauli set, and the process repeats $N$ times to get $N$ snapshots. (d) One ZECS application is quantification of the entanglement entropy from disjoint circuits. First, the full-density state operator $\rho_{T}$ is reconstructed, and then the partial trace to the disjoint circuits is applied. If there is no leakage of information, the entanglement entropy $S_{AB}\approx 0$.
@@
In recent years, the capabilities of quantum processing units (QPUs) have grown to the size of hundreds of qubits [1, 2]. In these devices, the noise is the most critical aspect, and, therefore, ways to characterize it at scale are needed. Usually, characterization methods show individual qubit behavior, 2-qubit gates, or how well the device operates altogether. Different methods have been proposed for the characterization of quantum hardware, e.g., quantum state tomography [3] (QST), randomized benchmarking (RB)[4], quantum volume (QV) [5], cross-entropy benchmarking (XEB)[6], algorithmic qubits (AQ)[7], or error per layered gate (EPLG)[8].
@@
From these techniques, one that recovers detailed information about a quantum system is QST. It searches through all the degrees of freedom of the density operator $\rho$ which for $n_{q}$ qubits scales as $2^{2n_{q}+1}$. In practice, common methods to calculate the QST of $\rho$ for an error $\varepsilon$ require at least $O(4^{n_{q}}/\varepsilon^{2})$ copies of $\rho$ [9, 10] . Therefore, the amount of resources needed to reconstruct the density state operator for a few qubits is impractical on real devices.
@@
In this work, we propose a new methodology to characterize QPUs performance and crosstalk, to this end, we developed zero-entropy classical shadow (ZECS), a methodology that uses classical shadows (CS) [11] as a subroutine to reconstruct the density state operator of small QPUs subsystems.
@@
CS tool has been used to get information on the expected values of quantum systems. For example, in [12], CS is used to predict the ground state properties of many-body physics problems combining CS with a polynomial-time machine learning algorithm. It has also been applied to predicting quantum Fisher information [13], quantum process tomography [14], and information scrambling [15]. The effects of noise on classical shadows have been analyzed in [16], and methods to mitigate these effects, including real QPU implementations, have been demonstrated in [17, 15, 14].
@@
Figure 1(a) shows the workflow of the ZECS methodology. First, classical snapshots of the density state operator are collected using random single-qubit Clifford circuits. Then, the density state operator $\rho_{cs}$ of these sections is reconstructed using the mean value of the different snapshots. Up to this point, the reconstruction is unstable, and $\rho_{cs}$ is not positive semidefinite, a needed condition of the density state operator. To correct this, ZECS is proposed here. This methodology projects $\rho_{cs}$ into the pure state of the eigenvector associated with the largest eigenvalue of $\rho_{cs}$. This strategy makes the density state operator generated, $\rho_{zecs}$, describe a positive semidefinite and unit trace matrix.
@@
Since the number of samples required to reconstruct density state operators is expected to grow exponentially with the number of qubits, as in other QST methods, ZECS is employed to gain insights into the behavior of small sections of quantum processing units (QPUs) even if the qubits involved in the reconstruction are spatially apart. Fig. 1(b) shows a section of ibm_brisbane layout, where 3 disjoint circuits are used (1,2), (3,4,5,6), and (7,8,9). The ZECS methodology is flexible enough to reconstruct the 3 disjoint sections or combinations of them. Fig. 1(c) shows the reconstruction of one of the disjoint sections, qubits (7,8,9). Using $\rho^{zecs}_{7,8,9}$, a diagnostic can be done in terms of, for instance, the fidelity (F) or trace distance (D).
@@
Experimentally, ZECS is used to make a diagnostic in terms of $F$, (or infidelity, $1-F$), $D$, and the entanglement entropy ($S_{ij}$; see Sec.A.1) on three IBM QPUs: ibm_lagos, ibm_brisbane, and ibm_fez with 7, 127, and 156 qubits, respectively. As a testing protocol, the EfficientSU2 circuit of qiskit is employed [18] and the QAOA algorithm in the case of ibm_fez. The EfficientSU2 circuit consists of parametric layers of single-qubit operations of the SU(2) set and CNOT entanglement gates. What is shown here is that with only 6,000 snapshots, the $\rho_{zecs}$ of 2, 3, and 4 qubits can be recovered. The different $\rho_{zecs}$ are then utilized to create a map of the fidelity and the entanglement entropy for the entire layout of the device. This information is used as a routing technique to select the best chain of 20 qubits for an optimization application. It is shown that this methodology significantly improves the performance of the application compared to the heaviest qiskit transpilation technique.
@@ II.1 Classical Shadow
Classical shadow is a method for reconstructing an approximate classical description of a quantum system using a small number of measurements [11]. To reconstruct an $n$-qubit quantum state $\rho$ using N snapshots, random unitary gates $U_{i}$ are applied to $\rho$
@@
This method is restricted here to the Pauli basis measurement (see Eq. (S44) in [11]) so that each snapshot is given by
@@ II.2 ZECS
where $\Sigma=\mathrm{diag}(|\lambda_{1}|,|\lambda_{2}|,...,|\lambda_{n}|)$ is a diagonal matrix with the non-negative singular values of $\rho_{cs}$ in decreasing order, These coincide with the absolute values of the eigenvalues because $\rho_{cs}$ is hermitian. $V=[|\psi_{1}\rangle,|\psi_{2}\rangle,...,|\psi_{n}\rangle]$ is a matrix with the (orthonormal) eigenvectors of $\rho_{\text{cs}}$. The zero-entropy (ZE) step consists of truncating the information used to reconstruct the density state operator to only that of the largest eigenvalue. This step is analogous to the rank reduction in QST [19, 20, 21] and can be seen as completely removing the entropy of $\rho_{cs}$. The eigenvector $|\psi_{1}\rangle$ associated with the largest singular value of $\rho_{cs}$, i.e., $|\lambda_{1}|$ is then used to reconstruct the density state operator by
@@
Note that by the well-known Mirsky generalization of the Eckart-Young theorem [22], the closest approximation of $\rho_{cs}$ by means of a rank-one hermitian operator $A$ with respect to a unitarily invariant operator norm is given by $A_{1}=\mathrm{V}\Sigma_{1}\mathrm{V}^{\dagger}=|\lambda_{1}|\,\rho_{\text{zecs}}$ where $\Sigma_{1}=\mathrm{diag}(|\lambda_{1}|,0,...,0)$, i.e.,
@@
Numerical evidence is provided in Sec.A.1 that shows the zero entropy strategy to recover information of a mixed state. We use a random perturbation on a 2-qubit Bell state following the methodology in [23] and show that if the perturbation is not large, one can always recover more information using this zero-entropy (ZE) methodology. Evidence is given in terms of the fidelity, F, the trace distance, D, and the concurrence, C, for a description of these metrics.
II.3 Experiments
@@
To test the ZECS methodology on real quantum hardware, the IBM Falcon (ibm_lagos), Eagle (ibm_brisbane), and Heron (ibm_fez) devices [24, 25] are used. These are superconductive quantum processors [26] with a Heavy-Hex layout [27]. Fig. 2 shows (a) the circuit used for the ibm_lagos experiment and (b) the ibm_lagos layout. The circuit used is an EfficientSU2 gate from qiskit. This is a parametrized circuit with $4nN_{q}$ parameters, where $n$ is the number of repetitions of the circuit and $N_{q}$ is the number of qubits. The parameters are randomly selected from a uniform distribution with values between 0 and $\pi/2$. 10,000 snapshots are employed for this circuit, measuring the 7 qubits involved.
@@
In the case of ibm_brisbane, 1 and 10 repetitions of the EfficientSU2 gate are applied in pairs of qubits throughout the device. The parameters are randomly selected for different pairs of qubits using the same distribution employed with ibm_lagos. 6,000 snapshots are utilized to measure all the qubits at the end of the protocol. In the case of ibm_fez, the 2-qubit systems execute an LR-QAOA [28]protocol with p=10. This circuit consists of 20 CZ gates and 10 Rx gates.
@@
Figure 2: (a) The EfficientSU2 circuit used for the characterization of ibm_lagos. The circuit is repeated $n$ layers, and the parameters for each layer, the $\theta_{i}^{n}$, are randomly chosen from a uniform distribution between 00 and $\pi/2$; (b) The layout of ibm_lagos where the vertices represent the qubits and the edges the physical connectivity between them.
II.4 Routing application

### v3

@@ Abstract
As quantum processing units (QPUs) scale toward hundreds of qubits, diagnosing noise-induced correlations (crosstalk) becomes critical for reliable quantum computation. In this work, we introduce Zero-Entropy Classical Shadows (ZECS), a diagnostic tool that uses information of a rank-one quantum state tomography (QST) reconstruction from classical shadow (CS) information to make a crosstalk diagnosis. We use ZECS on trapped ion and superconductive QPUs including ionq_forte (36 qubits), ibm_brisbane (127 qubits), and ibm_fez (156 qubits), using from 1,000 to 6,000 samples. With these samples, we use the ZECS to characterize crosstalk among disjoint qubit subsets across the full hardware. This information is then used to select low-crosstalk qubit subsets on ibm_fez for executing the Quantum Approximate Optimization Algorithm (QAOA) on a 20-qubit problem. Compared to the best qubit selection via Qiskit transpilation, our method improves solution quality by 10% and increases algorithmic coherence by 33%. ZECS offers a scalable and measurement-efficient approach to diagnosing crosstalk in large-scale QPUs.
@@
zero-entropy classical shadow, quantum state tomography, routing qubits, IBM Quantum, QPU crosstalk, IonQ Forte.
@@ I Introduction
Figure 1: (a) Shows the ZECS density state operator reconstruction workflow. (b) Shows the QPU layout with three sets of qubits chosen to implement some circuits. Vertices represent qubits and edges two-qubit native interactions of the QPU. (c) Shows the acquisition of the CS information to reconstruct the density state operator. After implementing the circuit, the measuring basis changes randomly from the Pauli set, and the process repeats NN times to get NN snapshots. (d) One ZECS application is the quantification of the entanglement entropy from disjoint circuits. First, the full-density state operator $\rho_{T}$ is reconstructed, and then the partial trace to the disjoint circuits is applied. If there is no leakage of information, the entanglement entropy $S_{AB}\approx 0$.
In recent years, the capabilities of quantum processing units (QPUs) have grown to the size of hundreds of qubits [Bluvstein2024, Kim2023]. At this scale, sources of noise only visible at large circuit sizes, such as crosstalk, are emerging. Crosstalk, understood here as a platform-independent phenomenon, refers to errors in the execution of quantum circuits that originate from interactions with the immediate neighborhood of the targeted qubits and beyond. These include correlated errors between distant qubits and gate errors induced by qubits not directly involved in the gate operations [Sarovar_2020]. Different studies have investigated the nature and characterization of crosstalk, including [Gambetta_2012, Piltz_2014, Sheldon_2016, Rudinger_2019, Sarovar_2020, Debroy_2020].
@@
Usually, characterization methods show how noise affects individual qubits, 2-qubit gates, or quantum circuits. Different methods have been proposed for the characterization of quantum hardware, e.g., quantum state tomography [DAriano2002] (QST), randomized benchmarking (RB)[Emerson2005], quantum volume (QV) [Moll2018], cross-entropy benchmarking (XEB)[Boixo2018], algorithmic qubits (AQ)[Chen2023], or error per layered gate (EPLG)[McKay2023].
@@
From these techniques, one that recovers detailed information about a quantum system is QST. It searches through all the degrees of freedom of the density operator $\rho$, which for $n_{q}$ qubits scales as $2^{2n_{q}+1}$. In practice, common methods to calculate the QST of $\rho$ for a sampling error $\varepsilon$ require at least $O(4^{n_{q}}/\varepsilon^{2})$ copies of $\rho$ [Haah2017, Gebhart2023] with a lower bound of $O(3^{n_{q}}/\varepsilon^{2})$ [chen2024optimaltradeoffentanglementcopy]. As a result, calculating the density operator becomes computationally impractical, even for systems with only a few qubits on real quantum hardware.
@@
In this work, we propose a new methodology to characterize QPUs performance and crosstalk. To this end, we developed zero-entropy classical shadow (ZECS), a methodology that uses classical shadows (CS) [Huang2020] as a subroutine to reconstruct the density state operator of small QPUs subsystems.
@@
CS tool has been used to get information on the expected values of quantum systems. For example, in [Huang2022], CS is used to predict the ground state properties of many-body physics problems, combining CS with a polynomial-time machine learning algorithm. It has also been applied to predicting quantum Fisher information [Rath2021], quantum process tomography [Levy2024], and information scrambling [McGinley2022]. The effects of noise on classical shadows have been analyzed in [Koh2022], and methods to mitigate these effects, including real QPU implementations, have been demonstrated in [Jnane2024, McGinley2022, Levy2024].
@@
Figure 1(a) shows the workflow of the ZECS methodology. First, classical snapshots of the density state operator are collected using random single-qubit Clifford circuits. Then, the density state operator $\rho_{cs}$ of these sections is reconstructed using the mean value of the different snapshots. Up to this point, the reconstruction is unstable, and $\rho_{cs}$ is not positive semidefinite, a needed condition of the density state operator. To correct this, we use the rank-one reconstruction of the density state operator, which we denominate Zero Entropy CS (ZECS). This methodology projects $\rho_{cs}$ into the pure state of the eigenvector associated with the largest eigenvalue of $\rho_{cs}$, artificially removing the entropy associated with it. This strategy makes the density state operator generated, $\rho_{zecs}$, to describe a positive semidefinite and unit trace matrix.
@@
Since the number of samples required to reconstruct density state operators is expected to grow exponentially with the number of qubits, as in other QST methods, ZECS is employed to gain insights into the behavior of small sections of quantum processing units (QPUs) even if the qubits involved in the reconstruction are spatially apart.
The benefit of ZECS comes from its reusability: a single batch of measurements can be reused to estimate an arbitrary density state operator of any set of qubits from the QPU, requiring at most $O(d\log^{2}d))$ random settings, where $d=2^{n_{q}}$, and assuming a rank-one (pure state) reconstruction [Gross_2010]. Therefore, for a trace distance error, $\varepsilon$, the number of samples required is of $O(n_{q}^{2}*2^{n_{q}}/\varepsilon^{2})$. In contrast, conventional QST requires $O(3^{n_{q}}/\varepsilon^{2})$ samples for each pair of qubits involved in the protocol. While ZECS provides less resolution to distinguish different noise processes, we show that it is well-suited for detecting crosstalk.
Figure 1(b) shows a section of ibm_brisbane layout, where 3 disjoint circuits are used (1,2), (3,4,5,6), and (7,8,9). The ZECS methodology is flexible enough to reconstruct the 3 disjoint sections or combinations of them. Fig. 1(c) shows the reconstruction of one of the disjoint sections, qubits (7,8,9). Using $\rho^{zecs}_{7,8,9}$, a diagnostic can be done in terms of, for instance, the fidelity (F) or trace distance (D).
@@
Experimentally, ZECS is used to make a diagnostic in terms of $F$, (or infidelity, 1F1-F), $D$, and the entanglement entropy ($S_{ij}$; see Sec.A.1) on three IBM QPUs: ibm_lagos, ibm_brisbane, and ibm_fez with 7, 127, and 156 qubits, respectively, and IonQ ionq_forte [Chen_2024]. As a testing protocol, the EfficientSU2 circuit of qiskit is employed [Javadi-Abhari2024], the QAOA algorithm in the case of ibm_fez, and random circuits. With only 1000 samples, the $\rho_{zecs}$ of 4 qubits can be recovered. The different $\rho_{zecs}$ are then utilized to create a map of the fidelity and the entanglement entropy for the entire layout of the device. This information is used as a routing technique to select the best chain of 20 qubits for an optimization application. It is shown that this methodology improves the performance of the application compared to the heaviest qiskit transpilation technique.
@@ II.1 Classical Shadow
Classical shadow is a method for reconstructing an approximate classical description of a quantum system using a small number of measurements [Huang2020]. To reconstruct an $n$-qubit quantum state $\rho$ using N snapshots, random unitary gates $U_{i}$ are applied to $\rho$
@@
This method is restricted here to the Pauli basis measurement (see Eq. (S44) in [Huang2020]) so that each snapshot is given by
@@ II.2 ZECS
where $\Sigma=\mathrm{diag}(|\lambda_{1}|,|\lambda_{2}|,...,|\lambda_{n}|)$ is a diagonal matrix with the non-negative singular values of $\rho_{cs}$ in decreasing order, These coincide with the absolute values of the eigenvalues because $\rho_{cs}$ is hermitian. $V=[|\psi_{1}\rangle,|\psi_{2}\rangle,...,|\psi_{n}\rangle]$ is a matrix with the (orthonormal) eigenvectors of $\rho_{\text{cs}}$. The zero-entropy (ZE) step consists of truncating the information used to reconstruct the density state operator to only that of the largest eigenvalue. This step is analogous to the rank reduction in QST [gross2010, Baldwin_2016, franca2021] and can be seen as completely removing the entropy of $\rho_{cs}$. The eigenvector $|\psi_{1}\rangle$ associated with the largest singular value of $\rho_{cs}$, i.e., $|\lambda_{1}|$ is then used to reconstruct the density state operator by
@@
Note that by the well-known Mirsky generalization of the Eckart-Young theorem [Golub1987], the closest approximation of $\rho_{cs}$ by means of a rank-one hermitian operator $A$ with respect to a unitarily invariant operator norm is given by $A_{1}=\mathrm{V}\Sigma_{1}\mathrm{V}^{\dagger}=|\lambda_{1}|\,\rho_{\text{zecs}}$ where $\Sigma_{1}=\mathrm{diag}(|\lambda_{1}|,0,...,0)$, i.e.,
@@
Numerical evidence is provided in Sec.A.1 that shows the zero entropy strategy to recover information of a mixed state. We use a random perturbation on a 2-qubit Bell state following the methodology in [Montanez-Barrera2022] and show that if the perturbation is not large, one can always recover more information using this zero-entropy (ZE) methodology. Evidence is given in terms of the fidelity, F, the trace distance, D, and the concurrence, C, for a description of these metrics.
II.3 Crosstalk characterization
Among the various techniques to quantify crosstalk, Simultaneous Randomized Benchmarking (SRB) [Gambetta_2012, mckay2020correlated] is used to assess whether running randomized tests in parallel reduces the performance of otherwise independent systems. SRB consists of three experiments: two involve performing RB on one system while the other remains idle, and the third runs RB on both systems simultaneously. This yields two error rates $r_{i}$ for the idle case and $r_{s}$ for the simultaneous case. The crosstalk is then quantified as $r_{i}-r_{s}$, representing the change in RB performance of each subsystem due to concurrent operation.
Another way to quantify crosstalk is through gate set tomography (GST), which can be used to determine whether the observed data is better explained by a crosstalk-free model, a model with context-dependent crosstalk between nearby qubits, general crosstalk, or non-Markovian noise. The most accurate representation is identified based on a metric called the likelihood [Rudinger_2021].
In this work, we use the entanglement entropy to characterize crosstalk. The entanglement entropy is expressed as
$S_{ab}=-\rho_{a}\log(\rho_{a})=-\rho_{b}\log(\rho_{b}),$ (10)
where $\rho_{a}={\rm Tr}_{b}(\rho)$ is the partial trace of the composite density state operator $\rho$ over the basis states of subsystem $b$. The classical cost to compute the entanglement entropy from a given state $\rho$ is dominated by two steps. The first entails forming the reduced density matrix $\rho_{a}=\operatorname{Tr}_{b}\rho$, which requires $O(d_{a}^{2}d_{b})$ arithmetic operations, where $d_{a}=2^{n_{q}^{a}}$ and $d_{b}=2^{n_{q}^{b}}$ are the Hilbert-space dimensions of subsystems $a$ and $b$. The second consists of obtaining the spectrum required by Eq. 10 (e.g., via SVD or diagonalization of $\rho_{a}$), which costs $O(d_{a}^{3})$.
When subsystems $a$ and $b$ are disentangled, i.e., $\rho=\rho_{a}\otimes\rho_{b}$, $S_{ab}=0$. In our experiments, we run quantum circuits on subsets of qubits independently, reconstructing $\rho$ using ZECS, so the entanglement entropy of combinations of these subsystems is expected to be zero. Uncorrelated noise should not affect the entanglement entropy. However, if crosstalk induces correlations between the subsystems, the entanglement entropy will deviate from zero
Figure 2 shows how ZECS reconstruction can be used to identify crosstalk in noisy quantum circuits. Fig. 2(a) depicts the circuit composed of two subsystems, $a$ and $b$. Crosstalk is simulated by introducing additional random two-qubit gates acting between one qubit of subsystem $a$ and one qubit of subsystem $b$, with the interaction strength controlled by the parameter $\varepsilon$. We consider random circuits with a two-qubit depth of 10, simulated using Qiskit’s noisy simulator [ibm_build_noise_models] with 2-qubit depolarizing noise $\lambda=1\times 10^{-3}$. At each depth step, a random two-qubit gate is applied to each subsystem, followed by a crosstalk gate. Three scenarios are investigated: $\varepsilon=0$ (no crosstalk), $\varepsilon=0.1$, and $\varepsilon=0.3$. The crosstalk gates are implemented as $R_{xx}(\theta)$ and $R_{zz}(\theta)$ rotations, with $\theta$ randomly chosen in the interval $\varepsilon,\varepsilon]$. As $\varepsilon$ increases, the effective strength of the induced crosstalk also increases.
Figure 2(b) shows how the entanglement entropy changes with the number of shadows for three different crosstalk strengths. Since depolarizing noise is included in all cases, the entanglement entropy approaches zero when $\varepsilon=0$, indicating that uncorrelated noise does not produce a large $S_{ab}$. In contrast, increasing the crosstalk strength consistently increases the observed $S_{ab}$, showing that ZECS is an effective method for detecting crosstalk. The inset of Fig. 2(b) presents the corresponding fidelity results, where stronger crosstalk also leads to lower fidelity. Together, these simulation outcomes clarify why ZECS serves as a useful methodology for identifying crosstalk.
Figure 2: Crosstalk simulation between subsystems $a$ and $b$. (a) A random circuit used to simulate crosstalk between subsystems $a$ and $b$. The noisy gate that represents the crosstalk connects both subsystems and has a strength parameter $\varepsilon$. (b) Entanglement entropy versus the number of shadows used for different crosstalk strengths for the simulation of the circuit in (a). The inset plot shows the fidelity for the same process.
II.4 Experiments
To evaluate the ZECS methodology on real quantum hardware, we use the IonQ Forte (ionq_forte) and the IBM Falcon (ibm_lagos), Eagle (ibm_brisbane), and Heron (ibm_fez) processors [IBMQuantumProcessors, IBM2021]. ionq_forte is a trapped ion device with a fully connected layout, while the other QPUs are superconductive quantum processors [Koch2007] with a Heavy-Hex layout [IBMHeavyHexLattice].
@@ II.4 Experiments
Fig. 3 shows (a) the circuit used for the ibm_lagos experiment and (b) the ibm_lagos layout. The circuit used is an EfficientSU2 gate from qiskit. This is a parametrized circuit with $4nN_{q}$ parameters, where $n$ is the number of repetitions of the circuit and $N_{q}$ is the number of qubits. The parameters are randomly selected from a uniform distribution with values between 0 and $\pi/2$. 10,000 snapshots are collected for this circuit on ibm_lagos, measuring the 7 qubits involved.
@@
In the case of ibm_brisbane, 1 and 10 repetitions of the EfficientSU2 gate are applied in pairs of qubits throughout the device. The parameters are randomly selected for different pairs of qubits using the same distribution employed with ibm_lagos. 6,000 snapshots are utilized to measure all the qubits at the end of the protocol. In the case of ibm_fez and ionq_forte, the 2-qubit systems execute an LR-QAOA [montanezbarrera2024universal] protocol with p=10. This circuit consists of 20 CZ gates and 10 Rx gates.
@@
Figure 3: (a) The EfficientSU2 circuit used for the characterization of ibm_lagos. The circuit is repeated $n$ layers, and the parameters for each layer, the $\theta_{i}^{n}$, are randomly chosen from a uniform distribution between 0 and $\pi/2$; (b) The layout of ibm_lagos where the vertices represent the qubits and the edges the physical connectivity between them.
II.5 Routing application